Can someone help me solve this step by step?
Simplify the expression.
[sqrt(12)-sqrt(2)][sqrt(12)+sqrt(2)]

Answer by rothauserc(4718)

[sqrt(12)-sqrt(2)][sqrt(12)+sqrt(2)]
let x = sqrt(12) and y = sqrt(2), then we have
(x -y) * (x +y)
x^2 -xy +xy -y^2
xy terms cancel
x^2 -y^2
12 -2 = 10